Establish the equation of the parabola with focus (-3,-4) and directrix 5x+6y-5=0
1
Expert's answer
2021-09-12T23:51:26-0400

A parabola is defined as follows:

For a given point, called the focus, and a given line not through the focus, called the directrix, a parabola is the locus of points such that the distance to the focus equals the distance to the directrix.

Let "P(x,y)" be the point on the parabola. Then


"(x+3)^2+(y+4)^2=\dfrac{(5x+6y-5)^2}{5^2+6^2}"

"61x^2+366x+549+61y^2+488y+976"

"=25x^2+36y^2+25+60xy-50x-60y"

"36x^2-60xy+25y^2+416x+548y+1500=0"

Therefore the required equation of the parabola is


"(6x-5y)^2+416x+548y+1500=0"